Kim has a car loan for $550 per month at 4% annual interest rate. She would like to pay the car off 2 years early. How much would the payoff amount be ?
Answer by jorel1380(3719)

550/month=6600/yr
6600/yr + 4%=6600(1+0.04)=6864
6864/yr * 2 years=$13,728 amount of payoff
☺☺☺☺